In any given meal. Exmaple 10 grams of fat, 20 grams of carbohydrate, and 5 grams of protein.

1. Calculate the % of energy of each energy nutrient in a given number of total kcalorie meal.
2. Calculate total kcalorie value of a meal given the total grams of energy nutrients.

I know you multiply carbs and proteins by 4 and lipids by 9, I am confused on how to know when to just multipy and when to divide by the total.

Here is what I think is correct please let me know if I am.

1. Calculate the % of energy of each energy nutrient in a given number of total kcalorie meal. multiply carbs and proteins by 4 and lipids by 9, then / by the total

2. Calculate total kcalorie value of a meal given the total grams of energy nutrients. multiply carbs and proteins by 4 and lipids by 9 then add together
Answer #1

Energy in 10 grams of fat =10 X 9= 90 KCal

Energy in 20 grams of carbohydrate = 20 X 4 = 80 Kcal

Energy in 5 gms of protein = 5x4= 20 Kcal  

Ans 2 Total energy in a meal = 90+80+20= 190 Kcal

-------------------------------------------------------------------------------------------------------------------------------------------

Ans 1

% energy of fats in the meal = (90x100 ) / 190=47.37%

% energy of carbohydrates in a meal = 80x100/190= 42.1 %

% energy of proteins in a meal = 20x100/190=10.52%